Express Delivery, Inc. claims that it will deliver packages in less than 48 hours. Office Equipment Company uses Express Delivery, Inc.'s package delivery service for delivering merchandise to its most valued customers. To check the validity of the claim made by Express Delivery, Office Equipment Company checked the delivery time for the last 37 order deliveries and learned that sample mean is 50 hours. Assume that population standard deviation of delivery time for all packages delivered by Express Delivery is 4.9 hours. What is the value of the test statistic?

Select one:

A. 2.18

B. 2.48

C. 2.68

D. -2.18

E. -2.48

F. None of the above is the correct answer
Answer #1

Solution,

The null and alternative hypothesis is ,

H0 :  \mu = 48

Ha : \mu < 48

\bar x = 37

\sigma = 4.9

n = 50

Test statistic = z =

= (\bar x - \mu ) / \sigma / \sqrt n

= (37 - 48 ) / 4.9 / \sqrt 50

Test statistic = z = -0.32

correct option is = F